OpenStudy (anonymous):

x=(1,2,3) Y=(0,1,1) the triangle with corners at (0,0,0), the tip of ((X*Y)/(Y*Y)) Y (with tail at (0,0,0) and the tip of X (with tail at (0,0,0)) will always give us a right triangle? Why?

OpenStudy (anonymous):

OZ={0, 5/2, 5/2} XZ=OZ-OX={-1, 1/2, -1/2} OZ * OX = 0

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

but when i calculate the number, oz*ox is not 0.?

OpenStudy (anonymous):

Sorry XZ * OZ =0

OpenStudy (anonymous):

Misprint

OpenStudy (anonymous):

xz *zo?

OpenStudy (anonymous):

yes or xz*oz

OpenStudy (anonymous):

thank you